OpenStudy (anonymous):

What is the value of x? 6x - 2(x + 4) = 12

OpenStudy (anonymous):

will medal and fan

OpenStudy (jennithemeani):

6x-2(x+4)=12 1. distribute the -2 6x-2x-8=12 2. Add like terms 4x-8=12 3. Isolate x add 8 to both sides 4x=20 4. Divide by 4 x=5

OpenStudy (anonymous):

thanks